# Noise estimation of SDE from a single data trajectory

### v1

@@ 3.4 Proof of Step 5
where $a(X_{r})=\frac{\sigma^{\prime}(X_{r})}{\sigma(X_{r})}$ and $b(X_{r})=\frac{2\psi^{t_{i},t_{i+1}}_{21}+\sigma(X_{r})\sigma^{\prime}(X_{r})^{2}+\sigma(X_{r})^{2}\sigma^{\prime\prime}(X_{r})}{2\sigma(X_{r})}$, for $r\in[t_{i},t_{i+1})$. Finally, if we know $\mu(X_{0})$, and equation (37) has a unique solution for each interval $t_{i},t_{i+1})$, then we can solve the ODE using the simple Euler scheme as follows
$\mu(X_{t_{i}})=\mu(X_{t_{i-1}})+\left[a(X_{t_{i-1}})\mu(X_{t_{i-1}})+b(X_{t_{i-1}})\right](X_{t_{i}}-X_{t_{i-1}}),\quad\forall\;i=1,2,\dots,n.$
Thus, the solution of the above ODE for each interval $t_{i},t_{i+1})$ can be used to construct estimators of the values of $\mu(X_{t_{i}})$ for $i=1,\cdots,N.$ The proof is thus finished. ∎
4 Numerical Experiments
4.1 Model description
In this section, we demonstrate the performance of our method for estimating the drift, diffusion function, and noise vector across several numerical examples. We assume the process $X_{t}$ satisfies the dynamics as given in Eq. (5). We simulate sample paths using a standard time-discretization scheme, such as the Euler–Maruyama. With access to the data we compute its quadratic variation $X]_{\pi,t_{i}}$ for each time index $t_{i}:0\leq i\leq n$ where $t_{0}=0,t_{n}=T$, $n=100000$, and $T=1$ of a chosen uniform partition $\pi$. The quadratic variation satisfies the ordinary differential equation $d[X]_{\pi,t}=(\sigma(X_{t}))^{2}dt$. On estimated $Y_{t}=[X]_{\pi,t}$, we solve the regression problem (18) to get the functional form of $(\sigma(X_{t}))^{2}$ and hence $\sigma(X_{t})$. The $L_{0}$ penalty $\|\Xi_{\sigma}\|_{0}$ in Eq. (18) induces sparsity, but since we focus only on estimating $\sigma(X_{t})$ accurately, no regularization coefficient is included. We know that under risk-neutral probability $\mathbb{Q}$, the dynamics of $X_{t}$ satisfy $dX_{t}=\sigma(X_{t})dB_{t}^{\mathbb{Q}}$, which can be used to get the $B_{t}^{\mathbb{Q}}$ vector assuming it starts at zero. Now we have access to the primary process $(t_{i},B_{t_{i}}^{\mathbb{Q}})$ for $0\leq i\leq n$. We extend the features by using the idea of the iterated Stratonovich integral of the primary process up to order $|t-s|^{2-\epsilon}$ i.e., the integrals corresponding to the following coefficients
$\Psi_{2}:=\{\psi_{1},\psi_{2},\psi_{12},\psi_{21},\psi_{22},\psi_{222}\}.$
We choose a sub-partition $\{j_{k}\}_{k=0}^{M},M<<N$ of the partition $\pi$ such that $t_{j_{0}}=t_{0},t_{j_{M}}=t_{N}$ with $M=1000$. On each sub-interval $t_{j_{k}},t_{j_{k+1}}]$ of this sub-partition $X$ is regressed against above features to estimate the values of coefficients in the $\Psi_{2}$ at $t=t_{j_{k}}$ for $k=0,\dots,M$. The ODE (24) is then solved to get the vector of $\mu(X_{t_{j_{k}}})$ for each $t_{j_{k}}$. Furthermore, we use the discretization scheme (25) to predict the noise increment vectors $(\Delta B_{t_{j_{0}}}^{\mathbb{P}},\Delta B_{t_{j_{1}}}^{\mathbb{P}},\cdots,\Delta B_{t_{j_{M}}}^{\mathbb{P}})$ hence the noise $B^{\mathbb{P}}$ by assuming it starts at zero. Finally, we use the SSISDE (28) to recover the symbolic form of drift $\mu(X_{t})$ and diffusion $\sigma(X_{t})$. We describe the model-selection procedure, i.e., $k$-fold time–series cross-validation, the CV error $\delta$, and the choice of the optimal sparsity level, before presenting the numerical examples.
4.1.1 Model selection via $k$-fold time–series cross-validation
We adopt an elastic-net penalty Hastie et al., (2009) in the penalized identification problem (28), namely
with $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ the usual vector norms. The penalty strength $\alpha>0$ and mixing parameter $\rho\in[0,1]$ are hyperparameters that control sparsity and shrinkage in $(\Xi_{\mu},\Xi_{\sigma})$ ($\rho=1$ recovers Lasso; $\rho=0$ Ridge). We select $(\alpha,\rho)$ by $k$-fold time–series cross-validation Stone, (1974) (no shuffling across time): let $\{A_{j}\}_{j=1}^{k}$ partition the index set $\{1,\dots,N-1\}$ into $k$ contiguous, non-overlapping validation blocks, and define $B_{j}:=\bigcup_{\ell\neq j}A_{\ell}$ as the complementary training indices. For each candidate pair $(\alpha,\rho)$ and each fold $j$, we fit (28) on $B_{j}$ to obtain $(\widehat{\Xi}_{\mu}^{(j)},\widehat{\Xi}_{\sigma}^{(j)})$, and compute the mean squared one-step prediction error on $A_{j}$:
The $k$-fold CV score and its standard error are
4.1.2 Choice of the final model
Let $(\alpha_{\star},\rho_{\star})=\arg\min_{\alpha,\rho}\overline{\delta}(\alpha,\rho)$ and define the one–standard-error Hastie et al., (2009) threshold
4.1.3 Optimally sparse size
where $\|\cdot\|_{0}$ counts nonzero entries [see e.g. Efron et al., (2004), Boninsegna et al., (2018)]. Grouping all grid points $(\alpha,\rho)$ that yield the same $n\in\mathbb{N}$, define
4.1.4 Practical settings used in our experiments
Unless stated otherwise, we use $k=7$ contiguous folds, a logarithmic grid $\Lambda=\{\alpha_{\min},\dots,\alpha_{\max}\}$ spanning several decades for $\alpha$, and $\mathcal{P}=\{0.3,0.4,0.5,0.7,0.85,0.95,1.0\}$ for $\rho$ (with $\rho=1.0$ recovering Lasso). Feature blocks are built as
with RMS column normalization applied separately to each block on the training portion of each fold. All reported coefficients are de-biased by restricted least squares on the selected support. Alongside $(\alpha^{\dagger},\rho^{\dagger})$ and $(\tilde{n}_{\mu},\tilde{n}_{\sigma})$, we provide the final model size, validation curves/surfaces, and $\delta$ vs. $n$ summaries to document the sparsity–accuracy trade-off achieved by (28) in each example. Finally, let us discuss the results of our numerical experiments in the case of different classical examples.
4.2 Black-Scholes model
Consider the Black-Scholes SDE
$dX_{t}=\mu\ X_{t}\ dt+\sigma\ X_{t}\ dB^{\mathbb{P}}_{t},$
where $\mu$ and $\sigma>0$ are two constants, and the solution is a Geometric Brownian motion. This is an example of an SDE that does not satisfy the ergodicity assumption. It is the widely used dynamics for stock prices in finance. The idea traces back to Osborne and Samuelson’s lognormal price hypothesis Osborne, (1959) and was formalized in the Black–Scholes–Merton framework Black and Scholes, (1973) as geometric Brownian motion. The drift term is linear in the price, $\mu X_{t}$, and shows a proportional (percentage) expected rate of return. The diffusion term is also linear, $\sigma X_{t}\,dB^{\mathbb{P}}_{t}$, and $\sigma$ is taken positive because it represents standard deviation, and any negative sign can be absorbed by flipping the Brownian motion. In short, the drift part governs the systematic trend while the diffusion part quantifies how volatile the price is around that trend.
For our numerical experiment we simulate $X_{t}$ by taking
$\mu(X_{t})=0.5X_{t},\qquad\text{and}\qquad\sigma(X_{t})=0.3X_{t}.$
To identify the Black–Scholes drift, diffusion function, and noise from a single path using the penalized residual (28). We take polynomial libraries for both blocks, $\Theta_{\mu}(x)=\{1,x,\ldots,x^{5}\}$ and $\Theta_{\sigma}(x)=\{1,x,\ldots,x^{5}\}$, and tune $(\alpha,\rho)$ by $k$-fold time–series CV with $k=7$ on a log grid for $\alpha$ and a discrete set for $\rho$, with de-biasing on the active set and the 1-SE rule for parsimony. The CV surface/curves select an Elastic–Net point
$\alpha^{\dagger}\approx 3.87\times 10^{-4},\qquad\rho^{\dagger}=0.85,$
lying in a broad near-optimal region (Fig. 1). At these hyperparameters, the recovered laws are
$\widehat{\mu}(x)=0.503681\,x,\qquad\widehat{\sigma}(x)=0.00241444+0.296498\,x,$
with one–step $\mathrm{MSE}=5.275\times 10^{-9}$. Thus, the drift is identified as purely proportional (support size $n_{\mu}=1$), and the diffusion is essentially linear with a small intercept (support size $n_{\sigma}=2$). The $\delta$–vs–$n$ summaries place the chosen sizes at the left edge of the flat accuracy region, indicating no material gain from adding higher-order terms (Fig. 2). The four-panel diagnostic shows that $\widehat{\mu}(X_{t})$ and $\widehat{\sigma}(X_{t})$ track the true trajectories closely, the recovered noise overlaps the driving process, and the simulated path is visually indistinguishable from the observed $X_{t}$ (Fig. 3). Overall, the identification is faithful to the Black–Scholes structure while remaining parsimonious.
Figure 1: Time–series CV over $(\alpha,\rho)$ with de-biasing on the active set; the selected point $(\alpha^{\dagger},\rho^{\dagger})=(3.87\times 10^{-4},\,0.85)$ lies on a broad near-optimal basin.
4.3 Quadratic coefficients SDE
In this example, we study the SDE with quadratic drift and diffusion,
which belongs to the broader class of polynomial/CEV–type diffusions Davydov and Linetsky, (2001). Unlike geometric Brownian motion, (38) features state–dependent nonlinearities in both the drift and the diffusion.
We illustrate the full identification pipeline on this one-dimensional SDE and recover both the drift and diffusion function from a single trajectory via the penalized residual (28). The feature libraries for drift and diffusion are polynomial dictionaries
The CV surface/curves in Fig. 4 show a broad near-optimal basin around $\alpha\sim 10^{-3}$ for $\rho\in[0.4,0.7]$. The 1-SE rule selects an Elastic–Net model at
$\alpha^{\dagger}\approx 8.73\times 10^{-4},\qquad\rho^{\dagger}=0.50,$
marked in the figure. This choice lies on the left edge of the plateau and yields a sparse, stable model with validation error statistically indistinguishable from the minimum.
Aggregating CV outcomes by support size $n=\|(\Xi_{\mu},\Xi_{\sigma})\|_{0}$ produces the characteristic $\delta(n)$ profile in Fig. 5: a sharp drop from the underfit regime, a flat region of near-minimal error, and a mild rise at larger supports. The vertical markers show the optimally sparse sizes $\tilde{n}$ returned by the 1-SE rule together with the sizes of the final full-data models. In our run, the chosen model sits at the left boundary of the flat region, indicating that additional nonzeros do not materially improve validation error, whereas removing one more term increases $\delta$ noticeably.
Finally, fitting (28) on the full dataset with $(\alpha^{\dagger},\rho^{\dagger})$ and de-biasing on the selected support yields
$\widehat{\mu}(x)=0.620512\,x+0.35905\,x^{2},\qquad\widehat{\sigma}(x)=-0.017782+0.535985\,x+0.0842435\,x^{2},$
with one–step MSE $1.448\times 10^{-7}$ (Fig. 6). The drift support matches the analytic basis $\{x,x^{2}\}$ and displays small coefficient biases. The diffusion includes a small intercept and mild biases on the linear and quadratic terms.
[…]
Figure 4: Time–series cross-validation over $(\alpha,\rho)$ with de-biasing on the active set. The selected point $(\alpha^{\dagger},\rho^{\dagger})=(8.73\times 10^{-4},\,0.50)$ lies on a broad plateau of optimal values.
Figure 5: Validation error $\delta(n)$ vs. support size $n$ with $\pm 1$ SE bands. Vertical lines indicate the optimally sparse sizes $\tilde{n}$ (1-SE rule) and the final full-data sizes.
Figure 6: (a) The true $\mu(X_{t})$ and reconstructed $\widehat{\mu}(X_{t})$ and (b) The true $\sigma(X_{t})$ and reconstructed $\widehat{\sigma}(X_{t})$ are plotted against time on the interval $[0,1]$. (c) The true noise, i.e., Brownian motion $B^{\mathbb{P}}$ that was used to simulate the actual path, the predicted noise, and the $\mathbb{Q}$-Brownian motion are plotted against time. (d) The actual path of the data $X_{t}$ and the reconstructed one from the $\widehat{\mu}(X_{t})$ and $\widehat{\sigma}(X_{t})$ are plotted against time. Each vector is of length $M$ with $M=1000$.

### v2

@@ Appendix D Proof of Step 5
Denote $a(X_{r})=\frac{\sigma^{\prime}(X_{r})}{\sigma(X_{r})}$ and $b(X_{r})=\frac{2\psi^{t_{i},t_{i+1}}_{21}+\sigma(X_{r})\sigma^{\prime}(X_{r})^{2}+\sigma(X_{r})^{2}\sigma^{\prime\prime}(X_{r})}{2\sigma(X_{r})}$, for $r\in[t_{i},t_{i+1})$. Since we know $\mu(X_{0})$, using Eular scheme for each interval $t_{i},t_{i+1})$, we can solve $\mu$ as follows
$\mu(X_{t_{i}})=\mu(X_{t_{i-1}})+\left[a(X_{t_{i-1}})\mu(X_{t_{i-1}})+b(X_{t_{i-1}})\right](X_{t_{i}}-X_{t_{i-1}}),\quad\forall\;i=1,2,\dots,n.$ (44)
Thus, we use estimate the values of $\mu(X_{t_{i}})$ for $i=1,\cdots,N.$ The proof is thus finished. ∎
Appendix E Proof of Proposition 5.1
The expansion of $X_{t}$ in Step 3 is a Taylor expansion for the following SDE,
$X_{t}=X_{s}+\int_{s}^{t}V_{0}(X_{r})dr+\int_{s}^{t}V_{1}(X_{r})\circ dB_{r}^{\mathbb{P}},$ (45)
where we denote the vector field $V_{0}(X_{r})=\tilde{\mu}(X_{r})$, and $V_{1}(X_{r})=\sigma(X_{r})$ to simplify the notation. According to [5, Theorem 1.6] (see also [28, Lemma 3.9] for a more general set-up of this expansion), for the one dimensional $X_{t}$, we have
$X_{t}=X_{s}+\sum_{k=1}^{N}g_{k}(t,s)+|t-s|^{\frac{N+1}{2}}R^{N+1}(t,s),$ (46)
where we set $\mathbf{B}^{0}_{t}=t$, and $\mathbf{B}^{1}_{t}=B^{\mathbb{P}}_{t}$, and introduce the following notation,
$g_{k}(t,s)=\sum_{|I|=k}P_{I}(X_{s})\int_{s<t_{1}<t_{2}\cdots<t_{k}<t}\circ\;d\mathbf{B}^{i_{1}}_{t_{1}}\cdots\circ d\mathbf{B}^{i_{k}}_{t_{k}},$ (47)
for a word $I=(i_{1},\cdots,i_{k})\in\{0,1\}^{k}$. Here we denote $|I|$ the size of $I$ which equals $k$, and and we denote $P_{I}(X_{s})=(V_{i_{1}}\cdots V_{i_{k}})(X_{s})$. In particular, $V_{i_{1}}\cdots V_{i_{k}}:=(V_{i_{1}}\cdots(V_{i_{k-2}}(V_{i_{k-1}}V_{i_{k}}))\cdots)$ represents the iterative operation of the vector fields in order. Following from [5, Proposition 4.3] (see also [28, Lemma 3.9] with $H=1/2$), there exists a random time $\zeta$, some constants $\alpha,c>0$, for every $\xi\geq 1$, such that
$\mathbb{P}\left(\sup_{t\in[t_{k},t_{k+1}]}\||t-t_{k}|^{\frac{N+1}{2}}R_{m+1}(t,t_{k})\|\geq\xi|t_{k+1}-t_{k}|^{\frac{N+1}{2}};\tau<\zeta\right)$
$\leq\exp\left(-\frac{c\xi^{\alpha}}{|t_{k+1}-t_{k}|}\right).$
The random time $\zeta$ is the exit time of $X_{t}$ in a compact domain $K$. We pick the compact domain $K$ large enough, such that $X_{t}$ stays inside $K$ for $t\in[t_{k},t_{k+1}]$. Since we take the Taylor expansion at level 3, by selecting $N=3$, we have
$\mathbb{P}\left(\sup_{t\in[t_{k},t_{k+1}]}\|R_{m+1}(t,t_{k})\|\geq\xi|t_{k+1}-t_{k}|^{2};\tau<\zeta\right)\leq\exp\left(-\frac{c\xi^{\alpha}}{|t_{k+1}-t_{k}|}\right).$
Applying the above estimate, denote $Z=\|X_{t}-L^{3}(t)\|^{2}$, where $L^{3}(t)=X_{t_{k}}+\sum_{k=1}^{3}g_{k}(t,t_{k})$ as defined in (46), we have
$\mathbb{E}[\|X_{t}-L^{3}(t)\|^{2}]=\mathbb{E}[Z]=\int_{0}^{\infty}\mathbb{P}(Z>z)dz\leq C|t_{k+1}-t_{k}|^{2},$
which completes the proof.
Appendix F Proof of Corollary 5.2
First, recall that the total approximation error of $X_{t}$ accumulated from Step 1-3 is coming from the error terms $\varepsilon^{\tilde{\mu}}$, $\varepsilon^{\sigma}$, $\varepsilon^{\tilde{\mu},\sigma}$ and $\varepsilon^{\mu/\sigma}$. For the Black-Scholes model since $\sigma(X_{t})=\sigma X_{t}$ and $\tilde{\mu}(X_{t})=\mu X_{t}-\sigma^{2}X_{t}$, the Equations (37) and (38) simplifies to
$\tilde{\mu}(X_{u})$ $=\tilde{\mu}(X_{s})+(\mu-\sigma^{2})\int_{s}^{u}\circ dX_{r}+0$
${\sigma}(X_{u})$ $=\sigma(X_{s})+\sigma^{2}\int_{s}^{u}\circ dX_{r}+0.$
This implies $\varepsilon^{\tilde{\mu}}=\varepsilon^{\sigma}=0$ in Step 2. Furthermore, $\frac{\mu(X_{s})}{\sigma(X_{s})}=\frac{\mu}{\sigma}$, we get $\varepsilon^{\mu/\sigma}=0$ in Step 4. Thus, we only need to estimate the error $X_{t}$ in Step 3, the proof follows directly from Proposition 5.1.